Suppose that your demand schedule for DVDs
is as follows:
Quantity Demanded Quantity Demanded
Price (income = $10,000) (income = $12,000)
$ 8 40 DVDs 50 DVDs
10 32 45
12 24 30
14 16 20
16 8 12
a. Use the midpoint method to calculate your
price elasticity of demand as the price of
DVDs increases from $8 to $10 if (i) your
income is $10,000 and (ii) your income is
$12,000.
b. Calculate your income elasticity of demand
as your income increases from $10,000 to
$12,000 if (i) the price is $12 and (ii) the price
is $16.
1
Expert's answer
2021-10-05T15:43:23-0400

Solution:

a.). Price elasticity of demand (PEd) = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


Using mid-point formula = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"

When income is 10,000:

Q2 = 32                  P2 = 10

Q1 = 40                  P1 = 8


PEd = "\frac{32 -40}{(32+40)/2 } \div \frac{10 -8}{(10+8)/2 } = \frac{ -8}{36 }\div \frac{ 2}{9 } = \frac{ -0.22}{0.22 } = -1"


PEd = 1, which is unit elastic.


When income is 12,000:

Q2 = 45                  P2 = 10

Q1 = 50                  P1 = 8


PEd = "\frac{45 -50}{(45+50)/2 } \div \frac{10 -8}{(10+8)/2 } = \frac{ -5}{47.5 }\div \frac{ 2}{9 } = \frac{ -0.11}{0.22 } = -0.48"


PEd = 0.48, which is unit inelastic.


b.). Income elasticity of demand (YEd) = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; Income}"


Using mid-point formula = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{I_{2} -I_{1}}{(I_{2}+I_{1})/2 }"

When Price is 12:

Q2 = 30                  I2 = 12,000

Q1 = 24                  I1 = 10,000


YEd = "\frac{30 -24}{(30+24)/2 } \div \frac{12,000 -10,000}{(12,000+10,000)/2 } = \frac{ 6}{27 }\div \frac{ 2000}{11,000 } = \frac{ 0.22}{0.18 } = 1.22"


Income elasticity of demand (YEd) = 1.22, which means that DVDs are income elastic. They are normal products and a luxury.

 

When Price is 16:

Q2 = 12                  I2 = 12,000

Q1 = 8                    I1 = 10,000


YEd = "\frac{12 -8}{(12+8)/2 } \div \frac{12,000 -10,000}{(12,000+10,000)/2 } = \frac{ 4}{27 }\div \frac{ 2000}{11,000 } = \frac{ 0.4}{0.18 } = 2.22"


Income elasticity of demand (YEd) = 2.22, which means that DVDs are highly income elastic. They are normal products and a luxury. As income increases, so is their demand.